A restaurant hands out a scratch-off game ticket with prizes being worth purchases at the restaurant. The back of the ticket lists the odds of winning each dollar value: 0.3 for $5, 0.005 for $30, 0.004 for $75, and 0.002 for $100. What are the odds that the ticket is worth at least $30?
Answer #1

probability that ticket is at least $30 =P(X=30)+P(X=75)+P(X=100)

=0.005/(1+0.005)+0.004/(1+0.004)+0.002/(1+0.002)

=0.010955

therefore odds that the ticket is worth at least $30 =probability/(1-probability)=0.010955/(1-0.010955)

=0.011077